OpenStudy (anonymous):

Find the slope of a line passing through (6, 7) and (2, 9)

OpenStudy (anonymous):

\[m=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} } \]\[m=\frac{ 9-7 }{ 2-6 }=\frac{ 2 }{ -4 } =\frac{ 1 }{ -2 }\]

OpenStudy (anonymous):

slope = change in y over change in x. Therefore m=(9-7)/(2-6) = -2/4 = -1/2

OpenStudy (anonymous):

Thanks (-:

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

you're welcome!